# thanks: Talento-CM fellow

### v1

@@ Abstract
With the advent of next-generation surveys, constraints on cosmological parameters are anticipated to become more stringent, particularly for the total neutrino mass. This study forecasts these constraints utilizing galaxy clusters from the Chinese Space Station Telescope (CSST). Employing Fisher matrix analysis, we derive the constraint $\sigma(M_{\nu})$ from cluster number counts, cluster power spectrum, and their combination. The investigation ranges from the standard cosmological model with neutrinos $\nu\Lambda$CDM to the inclusion of dynamic dark energy in the $\nu w_{0}w_{a}$ model, revealing a minor impact of dark energy on neutrino constraints. We examine the largest source of systematic uncertainty arising from the mass-observable relation and find that, with perfect knowledge of the scaling relation parameters, CSST clusters have the potential to enhance precision, tightening constraints to $0.034$ eV. We also study the effects of the maximum redshift $z_{max}$ and other uncertainties, including redshift, halo mass function, and bias. Furthermore, we emphasize the significance of accounting for the growth-induced scale-dependent bias (GISDB) effect, which is expected to have an impact of 1.5 to 2.2 times on the final constraint.
@@
In the standard model of particle physics, neutrinos are considered to be massless. However, a significant breakthrough in understanding neutrino properties occurred in 1998 when the Super-K atmospheric neutrino experiment in Japan confirmed the existence of oscillations [1], indicating that neutrinos can change from one flavor to another during their flight. This discovery provide evidence for non-zero neutrino masses, although they are extremely small.
Various oscillation experiments, including those involving solar, atmospheric, accelerator, and reactor neutrinos, have been conducted. These experiments can determine the differences in mass-squared values between the three different mass eigenstates of neutrinos. However, they cannot determine the absolute value of neutrino masses. As a result, neutrinos can be categorized into two possible scenarios: the normal hierarchy (NH), where the two lighter neutrinos are closer in mass, or the inverted hierarchy (IH), where the two heavier neutrinos are. The current neutrino oscillation experiment [2] measured:
@@
in which the subscript {1,2,3} represents different mass eigenstate. So the oscillation experiment can constrain the total neutrino mass $M_{\nu}\equiv\sum_{\nu=1,2,3}m_{\nu}>0.06$ eV for NH($m_{3}>m_{2}>m_{1}$), and $M_{\nu}>0.10$ eV for IH($m_{2}>m_{1}>m_{3}$).
@@
How to measure the absolute neutrino mass? There are three different approaches [3][4]: (i)neutrinoless double $\beta$-decay $(0\nu\beta\beta)$: the KamLAND-Zen experiment obtained $M_{\nu}<0.18-0.48$ eV at the 90% confidence level (C.L.) [5]; (ii)kinematic effects: the KATRIN tritium decay experiment obtained $M_{\nu}<1.1$ eV at 90% C.L. [6]. (iii)cosmology (see [7, 8] for reviews).
@@
During the early time of the Universe, neutrinos exhibit relativistic behavior, leading to a delay at the time of matter-radiation equality. Consequently, this delay suppresses the growth of perturbations since the majority of growth takes place during the matter-dominated era. Additionally, due to their small mass, neutrinos tend to freely stream out of overdense regions, which is known as the free streaming effect . This effect smears out fluctuations on scales smaller than the free-streaming scale $k_{\text{fs}}$ [7]:
$k_{\text{fs}}=\frac{0.82}{\sqrt{1+z}}\left(\frac{m_{\nu}}{1eV}\right){{\,h\,{\rm Mpc}^{-1}}}.$ (1)
@@
Considering the impacts of these factors on the expansion history of the Universe and the growth of cosmic structures, various techniques are employed to constrain the total neutrino mass, including the cosmic microwave background (CMB) primary anisotropies, CMB lensing, galaxy distribution, galaxy lensing, Lyman-$\alpha$ forest, 21 cm surveys, supernovae (see [9, 10] for reviews and their constraints). The current tightest constraint comes from combining observations of the Sloan Digital Sky Survey (SDSS) baryon acoustic oscillation (BAO) and redshift-space distortions (RSD), Planck CMB anisotropies, Pantheon Type Ia supernovae (SNe Ia), and Dark Energy Survey (DES) weak lensing and galaxy clustering measurements, resulting in an upper limit of $M_{\nu}<0.115$ eV at a 95% C.L. [11]. Many studies have been conducted to predict the constraints that can be achieved by next-generation surveys such as Large Synoptic Survey Telescope (LSST) or Vera C. Rubin Observatory [12], Euclid [13] and Square Kilometre Array (SKA) [14]. For instance, Yu et al. [15] find that a combination of LSST galaxies and CMB-S4 lensing can achieve constraints of $0.025$ eV, while Dey et al. [16] utilize 21-cm intensity mapping, galaxy clustering, cosmic shear observations from the SKA survey, and CMB primary anisotropies from CMB-S4 to obtain a constraint of $0.006$ eV. In this study, we focus on the probe of neutrino by using galaxy clusters [see 17, 18, for reviews] and provide a forecast for the Chinese Space Station Telescope (CSST), a next-generation survey[19, 20].
@@
As the largest gravitationally bound structures in the Universe, galaxy clusters, although less numerous than galaxies, have more accurate theoretical descriptions and smaller observed redshift errors. The typical size of a cluster is about $1{{\,h^{-1}{\rm Mpc}}}$, which is comparable to the $k_{\text{fs}}$. Therefore clusters are sensitive to neutrino mass. Specifically, the two most studied observables of clusters are: cluster number counts $N(z)$ and cluster power spectrum $P_{c}(k)$. These two are complementary[21]: neutrinos not only suppress the growth rate of perturbations in the linear regime, leading to a decrease in $N(z)$, but also affect the fluctuation on small scales, consequently influencing the shape of $P_{c}(k)$.
CSST is a 2-meter space telescope positioned on the same orbit as the China Space Station. It will cover an area of 17,500 deg2 on the sky over a period of ten years. Compared to other large-field surveys, CSST excels in depth and image resolution[22]. Another unique advantage is its capability to conduct seamless spectroscopic observations within the same sky regions. With these advantages, CSST is expected to provide more extensive and higher-quality data. As a result, the improved data from CSST is anticipated to enable more accurate constraints on the total neutrino mass.
@@
In this study, we aim to provide forecasts for constraints on the total neutrino mass $M_{\nu}$ using CSST galaxy clusters. The paper is organized as follows: In section II, we introduce our method, which is based on the Fisher matrix and incorporates neutrino prescriptions. Section III presents the main results of our analysis with different cosmological models and scenarios. In Section IV, we discuss factors that could impact our findings. Finally, we summarize and conclude our study in Section V.
@@
Throughout this paper, the natural logarithm is denoted by $\ln$, while $\log$ denotes the decimal logarithm. The halo mass is defined as $M\equiv M_{200m}$, i.e. cluster mass enclosed within a spherical overdensity of 200 with respect to the mean matter density, unless otherwise specified.
@@
In this section, we present the theoretical framework that describes cluster number counts and cluster power spectrum, as well as their Fisher matrix. Particularly, we perform some modifications due to the presence of massive neutrinos.
@@
Galaxy clusters are regarded as dark matter halos. The halo mass function (hereafter, HMF) is defined to be the comoving number density of clusters, denoted as $n(M,z)$, as a function of cluster mass $\ln M$ and redshift $z$. The HMF is given by the equation:
@@
Here, $f(\sigma)$ is a semi-analytical function calibrated against N-body simulations, which is related to the nonlinear collapse [23, 24, 25, 26, 27]. $\rho_{m}$ represents the matter density at the present day. The quantity $\sigma$ corresponds to the root mean square (rms) of the linear matter power spectrum $P_{m}$ and can be expressed as:
@@
where, $W(kR)$ represents the Fourier transform of the top-hat window function with a radius $R$, defined as
@@
The relationship between the halo mass $M$ and the radius $R$ in the top-hat window function is given by:
@@
The power spectrum $P_{m}$ in Eq.(3) is obtained from CLASS (Cosmic Linear Anisotropy Solving System) [28].
@@
where, $\Delta\Omega$ represents the solid angle covered by the survey, and $dV(z)/dzd\Omega$ denotes the comoving volume per unit redshift and solid angle.
Considering the actual survey, there is always an error between the true and observed redshift $[z,z^{ob}]$, as well as for cluster mass $[M,M^{ob}]$.
For the observed redshift, we employ a Gaussian distribution $P(z^{ob}\mid z)$ with an expected value of $z$ and a variance of $\sigma^{2}_{z^{ob}\mid z}$. This distribution represents the probability that the observed photometric redshift $z^{ob}$ corresponds to the true redshift $z$:
@@
where the dispersion of photometric redshift $\sigma_{z^{ob}\mid z}(z)$ generally increases with redshift in the form of $1+z$ [29], originating from poorer redshift estimates for faint galaxies at higher redshifts.
For the observed cluster mass, following [30] [31], we assume that the probability distribution function of the observed mass $M^{ob}$ for clusters with a fixed mass $M$ and redshift $z$ is given by a log-normal distribution with an expected value $\ln M^{ob}|\ln M,z\right>$, and a variance $\sigma^{2}_{\ln M^{ob}\mid\ln M,z}$:
@@
where $x(\ln M^{ob},\ln M,z)\equiv\ln M^{ob}-\left<\ln M^{ob}|\ln M,z\right>$ represents the deviation of the observed mass from its mean value. The mean value $\ln M^{ob}|\ln M,z\right>$ is given by:
$\ln M^{ob}|\ln M,z\right>=\ln M+\ln M_{\text{bias}},$ (10)
where $\ln M_{\text{bias}}$ is the mass bias, parameterized as [31]:
@@
$\ln\lambda^{ob}|M,z\right>$ $=\ln A_{\lambda}+B_{\lambda}\ln(\frac{M}{3\times 10^{14}M_{\odot}h^{-1}})$
$+C_{\lambda}\ln(\frac{1+z}{1+0.45}),$ (12)
Parameters $A_{\lambda},B_{\lambda}$, $C_{\lambda}$ and $D_{\lambda}$ here are also nuisance parameters. To estimate $\sigma_{\ln\lambda^{ob}}^{2}$, we use $D_{\lambda}$ in $\sigma_{\ln\lambda}$ to replace the constant term and add an additional term to represent the projection effects. For simplicity, we replace the Poisson term $1/\left<\lambda^{ob}(M)\right>$ with $1/\exp\left<\ln\lambda^{ob}(M)\right>$:
$\sigma_{\ln\lambda^{ob}}^{2}=D_{\lambda}^{2}+\frac{1}{\exp\left<\ln\lambda^{ob}(M)\right>}.$ (13)
Finally, the scatter in the observed mass can be expressed as:
@@
where the first term represents the scatter in the true mass $\sigma_{\ln M}$, derived from Eq.(II.1) and Eq.(13), and the second term accounts for an additional redshift-dependent projection effect, which is associated with the photometric redshift error and the algorithm employed by cluster finders [30]. The redshift coefficient $\kappa$ is a nuisance parameter.
It should be noted that, Costanzi et al. [36] use $M_{500c}$, defined as the cluster mass enclosed within a spherical overdensity of 500 with respect to the critical matter density. So, when applying Eq.(II.1), we convert mass from $M_{200m}$ to $M_{500c}$ assuming that the halo density profile is described by the Navarro-Frenk-White (NFW) model [38] with a concentration parameter from [39].
Let $\Delta\ln M^{ob}_{i}$ denotes the observed mass bin $i$, and $\Delta z_{j}^{ob}$ denotes the observed redshift bin $j$. Finally, the expectation value of the number of galaxy clusters $N_{ij}$ incorporating observed effects is given by:
@@
The cluster power spectrum can be calculated from the linear matter power spectrum $P_{m}$ [40]:
@@
with a bias $b_{\text{eff}}(z)$ and incorporating several effects, the linear redshift-space distortion (RSD) effect $\mathcal{R}(k^{\text{true}}_{\perp},k^{\text{true}}_{\|})$, Alcock-Paczynski (AP) effect $\mathcal{AP}(z)$, and redshift uncertainty $\mathcal{Z}(z)$. The wave numbers here are distinguished as fiducial $k^{\text{fid}}$ and true $k^{\text{true}}$, associated with the AP effect, while all wave numbers discussed in subsection II.1 are true $k=k^{\text{true}}$. Each term is detailed below.
The effective bias $b_{\text{eff}}(z)$ is averaged over all cluster bias $b(M,z)$ at redshift $z$:
@@
where, the mass observation effect $P(\ln M^{ob}|\ln M,z)$ has also been considered.
The RSD effect refers to the distortion of the observed clustering pattern of galaxies in redshift space due to their line-of-sight velocities, thus is closely related to the line-of-sight wave numbers $k^{\text{true}}_{\|}$:
@@
where $D_{\text{grow}}$ is the linear growth rate, and $a$ is the expansion factor normalized to unity today.
The AP effect arises from assuming a fiducial cosmology to infer the position of a galaxy from its observed redshift and angle. If the ’true’ background cosmology differs from the fiducial cosmology, the inferred comoving radial and transverse distances are distorted with respect to the true ones. We fix the fiducial wave numbers $k^{\text{fid}}$, then compute the true wave numbers $k^{\text{true}}$:
@@
The treatment for redshift uncertainty in the cluster power spectrum is different from that in the number counts Eq.(7). The cluster number distribution is ’smeared’ along the radial direction, damping the power spectrum by [41]:
@@
where $\sigma_{z^{ob}\mid z}(z)$ is the photometric error varying with redshift Eq.(8).

### v2

@@ Abstract
With the advent of next-generation surveys, constraints on cosmological parameters are anticipated to become more stringent, particularly for the total neutrino mass. This study forecasts such constraints utilizing galaxy clusters from the Chinese Space Station Telescope (CSST). Employing Fisher matrix techniques, we derive the constraint $\sigma(M_{\nu})$ from cluster number counts, cluster power spectrum, and their combination. The investigation covers both the standard cosmological model with massive neutrinos $\nu\Lambda$CDM and the inclusion of dynamic dark energy in the $\nu w_{0}w_{a}$ model, revealing a minor impact of dark energy on neutrino mass constraints. We examine the largest source of systematic arising from the mass-observable relation uncertainties and find that, with perfect knowledge of the scaling relation parameters, CSST clusters have the potential to enhance precision, tightening constraints to $\sim 0.03$ eV. We also study the effects of the maximum redshift $z_{max}$ and other uncertainties, including those in redshift, halo mass function, and bias. Furthermore, we emphasize the significance of accounting for the growth-induced scale-dependent bias (GISDB) effect, which we find can tighten the final constraint by a factor of $1.2$ $2.2$.
@@
In the standard model of particle physics, neutrinos are expected to be massless. However, a significant breakthrough in understanding neutrinos occurred in 1998 when the Super-Kamiokande atmospheric neutrino experiment in Japan confirmed neutrino oscillations [1], that is, neutrinos can change from one flavor to another during their propagation. This discovery provides evidence for non-zero neutrino masses, although small.
Various oscillation experiments, including those involving solar, atmospheric, accelerator and reactor neutrinos, have been conducted. They can determine the differences in squared masses between the three different mass eigenstates of neutrinos. However, they cannot determine the absolute value of neutrino masses. As a result, neutrinos can be organized into two possible scenarios: the normal hierarchy (NH), where the two lighter neutrinos are closer in mass, or the inverted hierarchy (IH), where the two heavier neutrinos are. The current neutrino oscillation experiments [2] measured:
@@
[…]
@@
How to measure the absolute neutrino mass? There are three different approaches [3][4]: (i) neutrinoless double $\beta$-decay $(0\nu\beta\beta)$ searches: the KamLAND-Zen experiment obtained the lightest neutrino mass $m_{\text{lightest}}<0.18-0.48$ eV at the 90% confidence level (C.L.) [5]; (ii) kinematic measurements: the KATRIN tritium decay experiment obtained $M_{\nu}<1.1$ eV at 90% C.L. [6]. (iii) cosmology (see [7, 8] for reviews).
@@
Fixing the matter density today, if neutrinos have mass, the non-relativistic component will decrease in the early Universe because neutrinos exhibit relativistic behavior at that time. This will lead to a delay in the time of matter-radiation equality.Consequently, this delay suppresses the growth of perturbations since the majority of growth takes place during the matter-dominated era. Additionally, due to their large thermal velocities, neutrinos tend to freely stream out of overdense regions, which is known as the free streaming effect. This effect tends to smear out fluctuations on scales smaller than the free-streaming scale $k_{\text{fs}}$ [7]:
$k_{\text{fs}}=0.82\frac{\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z)^{3}}}{(1+z)^{2}}\left(\frac{m_{\nu}}{1eV}\right){{\,h\,{\rm Mpc}^{-1}}},$ (1)
where $\Omega_{\Lambda}$ (resp. $\Omega_{m}$) is the cosmological constant (resp. matter) density fraction evaluated today.
@@
Considering the impacts of neutrinos on the expansion history of the Universe and the growth of cosmic structures, various probes are developed to constrain the total neutrino mass, including the cosmic microwave background (CMB) anisotropies, CMB lensing, galaxy clustering, galaxy lensing, Lyman-$\alpha$ forest, 21 cm intensity mapping (see [9, 10] for reviews). The current tightest constraint comes from combining baryon acoustic oscillation (BAO) data from the Dark Energy Spectroscopic Instrument (DESI), with CMB anisotropies and CMB lensing data from both Planck and Atacama Cosmology Telescope (ACT), resulting in an upper limit of $M_{\nu}<0.072(0.113)$ eV at a 95% C.L. for a $M_{\nu}>0(0.059)$ eV prior [11]. Many studies have been conducted to forecast constraints that can be achieved by next-generation surveys such as the Vera Rubin Observatory Legacy Survey of Space and Time (LSST) [12], Nancy Grace Roman Space Telescope [13] and Square Kilometre Array (SKA)[14]. For instance, Yu et al. [15] find that a combination of LSST galaxies and CMB-S4 lensing can achieve constraints on $M_{\nu}$ of $0.025$ eV. In this study, we focus on the probe of neutrino mass by using galaxy clusters [see 16, 17, for reviews] and provide a forecast for the Chinese Space Station Telescope (CSST), a next-generation space-based galaxy survey[18, 19].
@@
As the largest gravitationally bound structures in the Universe, galaxy clusters, although less abundant than galaxies, have more accurate theoretical descriptions and smaller observed redshift errors. Clusters are also sensitive to neutrino mass. Specifically, the two most studied observables of clusters are: cluster number counts $N(z)$ and cluster power spectrum $P_{c}(k)$. These two are complementary[20]: massive neutrinos not only suppress the overall amplitude of matter fluctuations, leading to a decrease in $N(z)$, but also affect the fluctuations on scales smaller than the free-streaming scales, consequently influencing the shape of $P_{c}(k)$.
CSST is a 2-meter space telescope operating in the same orbit as the China Manned Space Station. It will perform simultaneously a multi-band imaging and slitless spectroscopic survey with an area of 17,500 deg2 in about ten years. Compared to other large-field surveys, CSST excels in image quality, number of filter bands, and its unique coverage in the near ultraviolet etc [21]. Hence, CSST is expected to be highly competitive. It will also be complementary to other large projects of its time, such as Euclid [22] and the Roman Space Telescope [13]. As a result, we expect CSST can provide more accurate constraints on the total neutrino mass as well.
@@
In this work, we aim at forecasting constraints on the total neutrino mass $M_{\nu}$ by using CSST galaxy clusters, specifically the two cluster probes of cluster number counts and power spectrum. The rest of this paper is organized as follows: in section II, we introduce our methodology, which is based on the Fisher matrix technique, with careful incorporation of effects from massive neutrinos. Section III presents our main results for different models and assumptions. In Section IV, we discuss various effects that could impact our findings. Finally, we summarize our results and conclude in Section V.
@@
Throughout this paper, natural logarithm is denoted as $\ln$, while $\log$ stands for decimal logarithm. The halo mass is defined as $M\equiv M_{200m}$, i.e. cluster mass enclosed within a radius with an interior overdensity of 200 times the mean matter density, unless otherwise specified.
@@
In this section, we present our theoretical calculations for cluster number counts and cluster power spectrum, as well as their Fisher matrices. Particularly, we incorporate the effects of massive neutrinos on both observables.
@@ II.1 Cluster number counts
Galaxy clusters can be taken as dark matter halos, with their comoving number density $n$ described by the halo mass function (hereafter, HMF), denoted as $dn/d\ln M(M,z)$, as a function of cluster mass $\ln M$ and redshift $z$:
@@
Here, $f(\sigma)$ is a function calibrated against N-body simulations, which is related to the nonlinear collapse [23, 24, 25, 26, 27]. $\rho_{m}$ represents the matter density at the present day. The quantity $\sigma(R,z)$ is the amplitude of the linear matter fluctuations at redshift $z$, smoothed by a top hat window function with scale $R$, and can be expressed as:
@@
where, $W(kR)$ represents the Fourier transform of the top-hat window function with a radius $R$, given by
@@
The relationship between the halo mass $M$ and the scale $R$ in the top-hat window function is given by:
@@
The linear matter power spectrum $P_{m}$ in Eq.(3) is obtained from CLASS (Cosmic Linear Anisotropy Solving System) [28].
@@
where, $\Delta\Omega$ represents the solid angle covered by the survey, and $dV(z)/dzd\Omega$ denotes the comoving volume per unit redshift and solid angle intervals.
Considering an actual survey, there is always an error between the true and observed redshifts $[z,z^{ob}]$, as well as for cluster mass $[M,M^{ob}]$. For the observed redshift, we employ a Gaussian distribution for $z^{ob}$, $P(z^{ob}\mid z)$, with an expected value of $z$ and a variance of $\sigma^{2}_{z^{ob}\mid z}$:
@@
where the dispersion of observed redshift $\sigma_{z^{ob}\mid z}(z)$ generally increases with redshift with a factor of $1+z$ [29], reflecting poorer redshift estimates for faint galaxies at higher redshifts.
For the observed cluster mass, following [30] [31], we assume that the probability distribution function of the observed mass $M^{ob}$ for clusters with a fixed true mass $M$ and true redshift $z$ is given by a log-normal distribution with an expected value $\ln M^{ob}\right>(\ln M,z)$, and a variance $\sigma^{2}_{\ln M^{ob}\mid\ln M,z}$:
@@
where $x(\ln M^{ob},\ln M,z)\equiv\ln M^{ob}-\left<\ln M^{ob}\right>(\ln M,z)$ represents the deviation of the observed mass from its mean value. The mean value $\ln M^{ob}\right>(\ln M,z)$ is given by:
$\ln M^{ob}\right>(\ln M,z)=\ln M+\ln M_{\text{bias}},$ (10)
where $\ln M_{\text{bias}}$ is the bias in the logarithm of mass, parameterized as [31]:
@@
The parameters $B_{M,0}$ and $\alpha$ here will be simultaneously constrained with cosmological parameters, often referred to as the self-calibration scheme.
@@
$\ln\lambda^{ob}\right>(M,z)$ $=\ln A_{\lambda}+B_{\lambda}\ln(\frac{M}{3\times 10^{14}M_{\odot}h^{-1}})$
$+C_{\lambda}\ln(\frac{1+z}{1+0.45}).$ (12)
Without the projection effects, we would estimate $\sigma_{\ln\lambda^{ob}}^{2}$ also as the sum of a constant term, which we denote as $D_{\lambda}^{2}$ as well, and a Poisson term. That is,
$\sigma_{\ln\lambda^{ob}}^{2}=D_{\lambda}^{2}+\frac{1}{\exp\left<\ln\lambda^{ob}(M)\right>},$ (13)
which after scaled by the square of the scaling-relation slope $B_{\lambda}^{2}$, will lead to a scatter in the logarithm of mass. Finally, we model $\sigma_{\ln M^{ob}}^{2}$ as the direct sum of this scaled scatter and an extra term accounting for the projection effects:
@@
where the second term accounts for the redshift-dependent projection effect that is associated with redshift errors and the algorithm employed by the cluster finder [30].
It should be noted that, Costanzi et al. [36] use $M_{500c}$, defined as the cluster mass enclosed within a radius with interior mean density of 500 times the critical density. So, when applying Eq.(II.1), we convert mass from $M_{200m}$ to $M_{500c}$ assuming that the halo density profile is described by the Navarro-Frenk-White (NFW) model [38] with a concentration parameter from [39].
Let $\Delta\ln M^{ob}_{i}$ denote the $i$th observed mass bin, and $\Delta z_{j}^{ob}$ denote the $j$th observed redshift bin. Finally, the expectation value of the number of detected galaxy clusters $N_{ij}$ is given by:
@@ II.2 Cluster power spectrum
The cluster power spectrum $P_{c}$ can be calculated from the matter power spectrum $P_{m}$ [40]:
@@
with a bias factor $b_{\text{eff}}(z)$ and several other factors, accounting for the redshift-space distortion (RSD) effect $\mathcal{R}(k^{\text{true}}_{\perp},k^{\text{true}}_{\|})$, Alcock-Paczynski (AP) effect $\mathcal{AP}(z)$, and redshift uncertainty $\mathcal{Z}(z)$. The wave numbers here are distinguished as fiducial ones $k^{\text{fid}}$ and true ones $k^{\text{true}}$, due to the AP effect. Each term is detailed below.
The effective bias $b_{\text{eff}}(z)$ is an average of individual cluster bias $b(M,z)$ at redshift $z$:
@@
where the mass measurement uncertainty has also been considered.
The RSD effect refers to the distortion of the observed clustering pattern of clusters in redshift space due to their line-of-sight velocities, thus has an explicit dependence on the line-of-sight wave number $k^{\text{true}}_{\|}$:
@@
where $D_{\text{grow}}$ is the linear growth rate, and $a$ is the scale factor normalized to unity today.
The AP effect arises from assuming a fiducial cosmology to infer the position of a cluster from its observed redshift and angular position. If the ’true’ background cosmology differs from the fiducial cosmology, the inferred comoving radial and transverse distances are distorted with respect to the true ones. We fix the fiducial wave numbers $k^{\text{fid}}$, then compute the true wave numbers $k^{\text{true}}$:
@@
The treatment to account for redshift uncertainty in the cluster power spectrum is different from that in the number counts. The cluster distribution is ’smeared’ along the radial direction, which damps the power spectrum by [41]:
@@
where $\sigma_{z^{ob}\mid z}(z)$ is the cluster redshift uncertainty Eq.(8).